### v1

@@ 4 Real-world Dataset
basket 974 4 449 2 3 0 2 SAM [17]
drawer 884 4 452 2 2 1 0 SAM [17]
laptop 740 4 412 2 2 0 1 Mask-RCNN [9]
scissors 922 4 421 2 2 0 1 Mask-RCNN [9]
suitcase 813 4 381 2 2 0 1 Mask-RCNN [9]
@@
To evaluate the performance of OP-Align in real-world scenarios, we introduce our novel real-world dataset. The real-world dataset contains 5 categories of articulated objects, basket, laptop, suitcase, drawer, and scissors, captured by ASUS Xtion RGB-D camera. For each category, we randomly select 4 objects for training and 2 objects for testing. For each object, we set 8 random joint states and captured about 30 frames of RGB-D images for each. We also generated object segmentation masks predicted with detection models such as Mask-RCNN [9] or Segment Anything Model (SAM) [17]. The object point cloud can be generated by combining the depth channel of RGB-D images with a segmentation mask. Tab. 2 and Fig. 5 show an overview of this dataset. The annotation of the real-world dataset includes each part’s segmentation, rotation, and translation and each joint’s pivot and direction.
@@ 5 Experiments
Datasets: We use a synthetic dataset generated by authors of EAP [24] and our real-world dataset for evaluation. The synthetic dataset contains laptop, safe, oven, washer, and eyeglasses categories, selected from the mesh data in HOI4D [25] and Shape2Motion [40] dataset. We follow EAP [24]’s authors to render these mesh data into the partially observed point cloud, simulating the point cloud observation from a single-view camera.
@@
Baselines: For the synthetic dataset, we choose EAP [24] and 3DGCN [22] as self-supervised and supervised method baselines. We also report the results of a ICP algorithm, and NPCS [20] with EPN [3] backbone, which the authors of EAP [24] implemented. For the real-world dataset, we trained 3DGCN [22] and PointNet++ [33] as supervised method baselines.
@@
Evaluation Metrics: For the synthetic dataset, we follow EAP [24] and report the mean values of segmentation IoU, part rotation error, part translation error, joint direction error, and the distance from a point to a line as joint pivot error. For the real-world dataset, we follow category-level 6D object pose estimation methods [39, 6, 4] and choose the mean average precision (mAP) with multiple thresholds. An instance’s part pose is considered correct if the mean translation and rotation error of each part are both below the given thresholds. Specifically, we use thresholds $5,10,15$cm for translation, and $5^{\circ},10^{\circ},15^{\circ}$ for rotation. We also use the same thresholds for joint pivot and direction. For part segmentation, we use the mean value of intersection over union (IoU) of each part and thresholds of $75\%,50\%$ as metrics.
@@
Evaluation Strategies: Because OP-Align is a self-supervised model, it only predicts the relative poses of the input and the reconstruction instead of the poses defined by humans. Therefore, to evaluate our model’s performance, we need to determine the poses of the reconstruction parts. To achieve this, we follow EAP [24] and utilize ground truth labels from the training set. In preparation, for each training data and each part, a relative pose between the reconstruction and the input is obtained through Equation 6 by using a trained model, which, in combination with the ground truth pose, derives an estimated pose of each part of the reconstruction. We use these estimated poses to determine one common pose for each part of the reconstruction via a RANSAC-based method. For evaluation, we use the common pose as the pose of each part of the reconstruction. See supplement material for more details. We also note that for symmetric object categories such as basket, laptop, scissors and suitcase, the part segmentation is easily replaced with each other. For each object, among all possible permutations of indices of segmentation labels, we choose the permutation with the largest mean IoU over parts. The poses of parts are also permuted according to the chosen permutation.
@@
Training Settings: We trained a model for each category for $20{,}000$ iterations with a batch size of $24$. We used the Adam optimizer with a learning rate of $0.0001$ and halved the learning rate every $5{,}000$ iterations. The total loss is defined as $\lambda_{\mathrm{o}}\mathcal{L}_{\mathrm{o}}+\lambda_{\mathrm{p}}\mathcal{L}_{\mathrm{p}}+\lambda_{\mathrm{regS}}\mathcal{L}_{\mathrm{regS}}+\lambda_{\mathrm{regD}}\mathcal{L}_{\mathrm{regD}}+\lambda_{\mathrm{regW}}\mathcal{L}_{\mathrm{regW}}+\lambda_{\mathrm{regP}}\mathcal{L}_{\mathrm{regP}}+\lambda_{\mathrm{regA}}\mathcal{L}_{\mathrm{regA}}+\lambda_{\mathrm{regJ}}\mathcal{L}_{\mathrm{regJ}}$, where $(\lambda_{\mathrm{o}},\lambda_{\mathrm{p}},\lambda_{\mathrm{regS}},\lambda_{\mathrm{regD}},\lambda_{\mathrm{regW}},\lambda_{\mathrm{regP}},\lambda_{\mathrm{regA}},\lambda_{\mathrm{regJ}})=(10,10,100,10,10,10,10,10)$. We randomly sample 1024 points without RGB information from each object as input.
@@ 5.1 Results on the Synthetic Dataset
3DGCN [22] 94.05 11.61 0.093 0.084 9.78 - -
NPCS-EPN [20] - 11.05 0.080 0.147 15.20 - -
@@
We compare the performance of OP-Align on the partially observed point cloud from the synthetic dataset with other methods. As the results in Tab. 3, OP-Align exceeds other self-supervised methods by a large margin on multiple metrics. These results show that OP-Align can provide accurate joint and part pose prediction along with part segmentation. The visualization shown in Fig. 6 (left) demonstrates that object-level alignment can align the input with reconstruction holistically, and part-level alignment can align each part of the input with the corresponding part of the reconstruction. Also, thanks to the object-level alignment for reducing the pose variance, our method achieved higher part segmentation performance when compared with EAP [24]. However, the part segmentation performance still has room for improvement. Our assumption is that supervised 3DGCN [22] can directly learn segmentation with the geometric feature from the point cloud, while OP-Align leverages the difference of point distance between each part-level aligned input and the reconstruction for the indirect learning of segmentation probability with $\mathcal{L}_{\mathrm{p}}$. Especially in the region close to the joint, where points of part-level aligned inputs easily overlap, the point distance had no significant difference between each part-level aligned input, resulting in sub-optimal segmentation performance. We also compared our model in terms of inference speed and GPU memory with EAP [24]. OP-Align utilizes less GPU memory and achieves faster inference speed.
@@ 5.2 Results on the Real-world Dataset
3DGCN[22] 83.31 95.83 47.51 85.79 94.59 13.07 46.77 68.66
PointNet++[33] 19.83 42.20 21.06 57.38 75.56 4.47 23.25 39.82
@@
We conduct self-supervised training for OP-Align and compared the result with supervised 3DGCN [22] and PointNet++ [33] on the real-world dataset. The results are shown in Tab. 4 and Fig. 7 and the visualization is shown in Fig. 6 (right). OP-Align achieves results better than or comparable to PointNet++ [33] on all the metrics, and results comparable to 3DGCN [22] on part metrics, even without any annotations. However, similar to the results on the synthetic dataset, part segmentation learning with $\mathcal{L}_{\mathrm{p}}$ requires accurate point distance between part-level aligned inputs and the reconstruction, which is extremely challenging in real-world environments where outliers and missing points commonly exist. We also notice that our model still lags behind supervised methods in terms of the joint pivot and part translation metrics, as shown in laptop and suitcase visualization in Fig. 6. This phenomenon may be because the predicted joint pivots by our model, while capable of achieving part-level alignment, may not necessarily overlap with the actual joint pivots in reality. This also affects the performance of part translation based on joint movement.

### v2

@@ 4 Real-world Dataset
basket 974 4 449 2 3 0 2 SAM [sam]
drawer 884 4 452 2 2 1 0 SAM [sam]
laptop 740 4 412 2 2 0 1 Mask-RCNN [MaskRCNN]
scissors 922 4 421 2 2 0 1 Mask-RCNN [MaskRCNN]
suitcase 813 4 381 2 2 0 1 Mask-RCNN [MaskRCNN]
@@
To evaluate the performance of OP-Align in real-world scenarios, we introduce our novel real-world dataset. The real-world dataset contains 5 categories of articulated objects, basket, laptop, suitcase, drawer, and scissors, captured by ASUS Xtion RGB-D camera. For each category, we randomly select 4 objects for training and 2 objects for testing. For each object, we set 8 random joint states and captured about 30 frames of RGB-D images for each. We also generated object segmentation masks predicted with detection models such as Mask-RCNN [MaskRCNN] or Segment Anything Model (SAM) [sam]. The object point cloud can be generated by combining the depth channel of RGB-D images with a segmentation mask. Tab. 2 and Fig. 5 show an overview of this dataset. The annotation of the real-world dataset includes each part’s segmentation, rotation, and translation and each joint’s pivot and direction.
@@ 5 Experiments
Datasets: We use a synthetic dataset generated by authors of EAP [PLSEE] and our real-world dataset for evaluation. The synthetic dataset contains laptop, safe, oven, washer, and eyeglasses categories, selected from the mesh data in HOI4D [hoi4d] and Shape2Motion [shape2motion] dataset. We follow EAP [PLSEE]’s authors to render these mesh data into the partially observed point cloud, simulating the point cloud observation from a single-view camera.
@@
Baselines: For the synthetic dataset, we choose EAP [PLSEE] and 3DGCN [3DGCN] as self-supervised and supervised method baselines. We also report the results of a ICP algorithm, and NPCS [ANCSH] with EPN [EPN] backbone, which the authors of EAP [PLSEE] implemented. For the real-world dataset, we trained 3DGCN [3DGCN] and PointNet++ [pointnet++] as supervised method baselines.
@@
Evaluation Metrics: For the synthetic dataset, we follow EAP [PLSEE] and report the mean values of segmentation IoU, part rotation error, part translation error, joint direction error, and the distance from a point to a line as joint pivot error. For the real-world dataset, we follow category-level 6D object pose estimation methods [NOCS, GPV-pose, FS-net] and choose the mean average precision (mAP) with multiple thresholds. An instance’s part pose is considered correct if the mean translation and rotation error of each part are both below the given thresholds. Specifically, we use thresholds 5,10,155,10,15cm for translation, and $5^{\circ},10^{\circ},15^{\circ}$ for rotation. We also use the same thresholds for joint pivot and direction. For part segmentation, we use the mean value of intersection over union (IoU) of each part and thresholds of $75\%,50\%$ as metrics.
@@
Evaluation Strategies: Because OP-Align is a self-supervised model, it only predicts the relative poses of the input and the reconstruction instead of the poses defined by humans. Therefore, to evaluate our model’s performance, we need to determine the poses of the reconstruction parts. To achieve this, we follow EAP [PLSEE] and utilize ground truth labels from the training set. In preparation, for each training data and each part, a relative pose between the reconstruction and the input is obtained through Equation 6 by using a trained model, which, in combination with the ground truth pose, derives an estimated pose of each part of the reconstruction. We use these estimated poses to determine one common pose for each part of the reconstruction via a RANSAC-based method. For evaluation, we use the common pose as the pose of each part of the reconstruction. See supplement material for more details. We also note that for symmetric object categories such as basket, laptop, scissors and suitcase, the part segmentation is easily replaced with each other. For each object, among all possible permutations of indices of segmentation labels, we choose the permutation with the largest mean IoU over parts. The poses of parts are also permuted according to the chosen permutation.
@@
Training Settings: We trained a model for each category for 20,00020{,}000 iterations with a batch size of $24$. We used the Adam optimizer with a learning rate of $0.0001$ and halved the learning rate every 5,0005{,}000 iterations. The total loss is defined as $\lambda_{\mathrm{o}}\mathcal{L}_{\mathrm{o}}+\lambda_{\mathrm{p}}\mathcal{L}_{\mathrm{p}}+\lambda_{\mathrm{regS}}\mathcal{L}_{\mathrm{regS}}+\lambda_{\mathrm{regD}}\mathcal{L}_{\mathrm{regD}}+\lambda_{\mathrm{regW}}\mathcal{L}_{\mathrm{regW}}+\lambda_{\mathrm{regP}}\mathcal{L}_{\mathrm{regP}}+\lambda_{\mathrm{regA}}\mathcal{L}_{\mathrm{regA}}+\lambda_{\mathrm{regJ}}\mathcal{L}_{\mathrm{regJ}}$, where $(\lambda_{\mathrm{o}},\lambda_{\mathrm{p}},\lambda_{\mathrm{regS}},\lambda_{\mathrm{regD}},\lambda_{\mathrm{regW}},\lambda_{\mathrm{regP}},\lambda_{\mathrm{regA}},\lambda_{\mathrm{regJ}})=(10,10,100,10,10,10,10,10)$. We randomly sample 1024 points without RGB information from each object as input.
@@ 5.1 Results on the Synthetic Dataset
3DGCN [3DGCN] 94.05 11.61 0.093 0.084 9.78 - -
NPCS-EPN [ANCSH] - 11.05 0.080 0.147 15.20 - -
@@
We compare the performance of OP-Align on the partially observed point cloud from the synthetic dataset with other methods. As the results in Tab. 3, OP-Align exceeds other self-supervised methods by a large margin on multiple metrics. These results show that OP-Align can provide accurate joint and part pose prediction along with part segmentation. The visualization shown in Fig. 6 (left) demonstrates that object-level alignment can align the input with reconstruction holistically, and part-level alignment can align each part of the input with the corresponding part of the reconstruction. Also, thanks to the object-level alignment for reducing the pose variance, our method achieved higher part segmentation performance when compared with EAP [PLSEE]. However, the part segmentation performance still has room for improvement. Our assumption is that supervised 3DGCN [3DGCN] can directly learn segmentation with the geometric feature from the point cloud, while OP-Align leverages the difference of point distance between each part-level aligned input and the reconstruction for the indirect learning of segmentation probability with $\mathcal{L}_{\mathrm{p}}$. Especially in the region close to the joint, where points of part-level aligned inputs easily overlap, the point distance had no significant difference between each part-level aligned input, resulting in sub-optimal segmentation performance. We also compared our model in terms of inference speed and GPU memory with EAP [PLSEE]. OP-Align utilizes less GPU memory and achieves faster inference speed.
@@ 5.2 Results on the Real-world Dataset
3DGCN[3DGCN] 83.31 95.83 47.51 85.79 94.59 13.07 46.77 68.66
PointNet++[pointnet++] 19.83 42.20 21.06 57.38 75.56 4.47 23.25 39.82
@@
We conduct self-supervised training for OP-Align and compared the result with supervised 3DGCN [3DGCN] and PointNet++ [pointnet++] on the real-world dataset. The results are shown in Tab. 4 and Fig. 7 and the visualization is shown in Fig. 6 (right). OP-Align achieves results better than or comparable to PointNet++ [pointnet++] on all the metrics, and results comparable to 3DGCN [3DGCN] on part metrics, even without any annotations. However, similar to the results on the synthetic dataset, part segmentation learning with $\mathcal{L}_{\mathrm{p}}$ requires accurate point distance between part-level aligned inputs and the reconstruction, which is extremely challenging in real-world environments where outliers and missing points commonly exist. We also notice that our model still lags behind supervised methods in terms of the joint pivot and part translation metrics, as shown in laptop and suitcase visualization in Fig. 6. This phenomenon may be because the predicted joint pivots by our model, while capable of achieving part-level alignment, may not necessarily overlap with the actual joint pivots in reality. This also affects the performance of part translation based on joint movement.